# Perfect Matching in Product Graphs and in their Random Subgraphs

### v2

@@ 3.2 Typical properties of obstructions
Whp there are no obstructions such that one (or more) of the following holds,
  1. (a)
    $\ell_{1}\geq 5u$,
  2. (b)
    $s\geq 6u$,
  3. (c)
    $\ell_{3}\geq\frac{6u}{d^{2}}$ and $u\geq d^{2}$.
@@
We first claim that in any of the above cases, we would have that
$\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq 5du.$
Indeed, if $\ell_{1}=|V_{1}|\geq 5u$, since the graph $G$ is $d$-regular, $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq d\ell_{1}\geq 5du$. Similarly, since by Theorem 1.2, every component $K_{i}$ in $G_{p}[S]$ satisfies $e(K_{i},K_{i}^{C})\geq|K_{i}|(d-2(C-1)\log_{2}d)$, if $s\geq 6u$ then $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq(1-o(1))ds\geq 5du$. Finally, if $\ell_{3}\geq\frac{6u}{d^{2}}$, then, by Theorem 1.2 and by convexity arguments similar to Lemma 3.4, we have that
$\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})$ $>(\ell_{3}-1)d^{2}(d-2(C-1)\log_{2}d)$
$\geq 6du-d^{3}+2(C-1)\log_{2}d\geq 5du,$
where we used the assumption that $u\geq d^{2}$.
Fix $u$ and recall that $U$ has at most $du$ edges touching it. Note that any edge between two components $K_{i},K_{j}$ for $i,j\in[\ell]$ would contradict the fact that we consider an obstruction. Hence, the probability of an obstruction with $\sum_{i\in[\ell]}e(K_{i},K_{i}^{C})\geq 5du$ is at most
$\binom{n}{u}(1-p)^{5du-du}\leq n^{u}n^{-(1-\epsilon)4u}\leq n^{-\frac{5u}{2}},$
where we used $(1-p)^{d}=n^{-(1-\epsilon)}$, and recalling that choosing $U$ determines the obstruction. A union bound over all possible values of $u$, and the at most $n^{2}$ choices for $s$ and $\ell_{3}$ yields that the probability of an event violating the statement of the claim is at most
$3n^{2}\sum_{u=1}^{n}n^{-\frac{5u}{2}}\leq 3n^{2}\cdot 2n^{-\frac{5}{2}}=o(1).$
We are now ready to describe for some range of $u$ a set $M$ containing $U$ with relatively few components.
@@
Whp for every obstruction with $u\geq 2^{\frac{d}{10C}}$, there is a $M\subseteq V$ satisfying the following,
@@
    $G[M]$ contains at most $u/2^{\frac{d}{(100C)^{3}}}$ components.
@@
Suppose that $U$ is an obstruction with $u=|U|$, $|V_{1}|=\ell_{1}$ and $s=|S|$. By Claim 3.6, we have that whp there is no obstruction with $\ell_{1}\geq 4u$, as well as no obstruction with $s\geq 5u$, and no obstruction with $\ell_{3}\geq 10u/d^{2}$. We may thus assume that $\ell_{1}<4u,s<5u$ and $\ell_{3}<10u/d^{2}$. In particular, $\ell_{1}+\ell_{2}\geq u+1-10u/d^{2}$.
For every component $K_{i}$ in $G_{p}[S\cup V_{1}]$, we have that $E_{G_{p}}(K_{i},K_{i}^{C})\subseteq E(K_{i},U)$. Now, by Theorem 1.2,
$\sum_{K_{i}\in G_{p}[S\cup V_{i}]}e(K_{i},K_{i}^{C})\geq d\ell_{1}+(1+o(1))ds\geq d(\ell_{1}+\ell_{2})\geq d(u+1-10u/d^{2}),$
where in the second inequality we used that $s\geq 3\ell_{2}$. Thus, the probability that there is such an obstruction with $e(U,V_{1}\cup S)\leq\frac{du}{100C^{2}}$ is at most
$\binom{n}{u}(1-p)^{d(u+1-10u/d^{2}-\frac{u}{100C^{2}})}$ $\leq\left(\frac{en}{u}\right)^{u}n^{-\left(1-\frac{1}{95C^{2}}\right)u}\leq\exp\left\{\frac{u}{95C^{2}}\ln n+u-u\ln u\right\}$
$\leq\exp\left\{u\left(\frac{\ln n}{95C^{2}}-\frac{\ln 2\cdot d}{10C}\right)\right\}=o\left(n^{-3}\right),$
where we used $n\leq C^{d}$, $C\geq 2$ and $u\geq 2^{\frac{d}{10C}}$. Hence, by the union bound over the at most $n^{3}$ possible choices for $u$, $s$ and $\ell_{3}$, we have that whp any obstruction $U$ with $u\geq 2^{\frac{d}{10C}}$ has $\ell_{1}<4u$, $s<5u$ and there are at least $\frac{du}{100C^{2}}$ edges in $G[U\cup V_{1}\cup S]$.
Let $M=U\cup V_{1}\cup S$. We have that $U\subseteq M$ and $|M|=u+\ell_{1}+s\leq 10u$, by our assumptions $\ell_{1}<4u$ and $s<5u$. 1.2 implies that for any set $M\subseteq V(G)$
@@
If $G[M]$ had at least $u/2^{\frac{d}{(100C)^{3}}}$ components, then by standard convexity arguments, we would have that
$e_{G}(M)$ $\leq|M|(C-1)\log_{2}\left(\frac{|M|}{u/2^{\frac{d}{(100C)^{3}}}}\right)$
$\leq 10u(C-1)\log_{2}\left(10\cdot 2^{\frac{d}{(100C)^{3}}}\right)\leq\frac{du}{(100C)^{2}}<\frac{du}{100C^{2}},$
@@
3.3 Typically no obstructions
We are now ready to show that whp there are no obstructions with $u\geq 1$ in $G_{p}$. We consider several cases separately: when $|U|$ is small; when $s=b=0$; when $s\neq 0$ and $b=0$; when $0\neq b\leq\frac{n}{2}$; and, when $b>\frac{n}{2}$. We show that for each of these cases, whp there are no such obstructions. Note that all events we consider happen with high probability, and thus their intersection also happens with high probability, thus completing the proof of Theorem 3b.
Obstructions with small $U$.
We denote by $\mathcal{B}_{u}$ the event there exists an obstruction in $G_{p}$ with $|U|=u$.
@@ 3.3 Typically no obstructions
We have that
$\bigcup_{1\leq u\leq\frac{n}{d^{C^{3}/p}}}\mathbb{P}\left(\mathcal{B}_{u}\right)=o(1).$
@@
Suppose there exists a construction with $|U|=u$ and $u\in\left[\frac{n}{d^{C^{3}/p}}\right]$. By Lemma 3.4, we have that
$\sum_{j=1}^{j=w/2}e(W_{j},W_{j}^{C})+\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})$ $\geq w(d-1)+d(\ell-1)+f(k-\ell+1)$
$\geq w(d-1)+du+f(n-2u-w)\geq du+f(2u),$
where we used that $\ell\geq u+1,k=n-u-w$ and the definition of $f(x)$.
We continue by considering two ranges separately. First, suppose that $u\leq 2^{\frac{d}{10C}}$. We then have that $(C-1)\log_{2}u\leq(C-1)\frac{d}{10C}<\frac{d}{10}$ and thus $f(2u)\geq 2u(d-(C-1)\log_{2}u)>\frac{3du}{2}$. As there are at most $du$ edges touching $U$, and noting that any edge between any two different components in $G_{p}[V\setminus U]$ would rule out the existence of such an obstruction, we have that
$\mathbb{P}\left(\mathcal{B}_{u}\right)\leq\binom{n}{u}(1-p)^{3du/2}\leq n^{-(1-\epsilon)\frac{3u}{2}+u}<n^{-u/3}.$
Therefore, by the union bound,
$\bigcup_{1\leq u\leq 2^{\frac{d}{10C}}}\mathbb{P}\left(\mathcal{B}_{u}\right)\leq\sum_{u=1}^{2^{\frac{d}{10C}}}\frac{1}{n^{u/3}}=o(1).$
@@
We now turn to $2^{\frac{d}{10C}}\leq u\leq\frac{n}{d^{C^{3}/p}}$. By Lemma 3.7, we may assume that there exists $M$ such that $|M|\leq 10u$, $U\subseteq M$ and there are at most $u/2^{\frac{d}{(100C)^{3}}}\eqqcolon r$ components in $G[M]$. We have at most $n^{r}$ ways to choose the sizes of these components. Thus, by Lemma 1.3, there are at most $\binom{10u}{u}n^{2r}(ed)^{10u}$ ways to choose $U$. Since $u\leq\frac{n}{d^{C^{3}/p}}$, we have that
$f(2u)\geq\frac{2u(d-\log_{C}u)}{C}\geq\frac{2uC^{2}\log_{C}d}{p}.$
@@
Thus, the probability of such an obstruction is at most
$\binom{10u}{u}n^{2r}(ed)^{10u}(1-p)^{\frac{2uC^{2}\log_{C}d}{p}}$ $\leq\left(10e\right)^{u}n^{2r}(ed)^{10u}\exp\left\{-2uC^{2}\log_{C}d\right\}$
$\leq\exp\left\{u\left(11+\ln 10+\frac{2\ln n}{2^{d/(100C)^{3}}}+10\ln d-2C^{2}\log_{C}d\right)\right\}$
$\leq\exp\left\{u\left(15+\frac{2\ln n}{2^{d/(100C)^{3}}}-\left(10-\frac{2C^{2}}{\ln C}\right)\ln d\right)\right\}$
$=\exp\left\{-\frac{u\ln d}{2}\right\}\leq d^{-u/2},$

### v3

@@ 3.3 Proof of Theorem 3b
Fix an obstruction. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components of $G_{p}[S]$. Then,
$\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq 3(d-1)(\ell_{2}-1)+f^{*}\left(s-3(\ell_{2}-1)\right).$
@@
By definition, $\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq\sum_{i=1}^{\ell_{2}}f^{*}(|S_{i}|)$. Further, we have that $\sum_{i=1}^{\ell_{2}}|S_{i}|=s$, and that for each $i\in[\ell_{2}]$, $3\leq|S_{i}|<n/d^{C^{3}/p}$. Thus, by Remark 3.6 and by standard convexity arguments, we have that $\sum_{i=1}^{\ell_{2}}f^{*}(|S_{i}|)$ is minimised when $|S_{i}|=3$ for all $i\in[1,\ell_{2}-1]$, and $|S_{\ell_{2}}|=s-3(\ell_{2}-1)$. Therefore,
$\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})$ $\geq 3(d-1)(\ell_{2}-1)+f^{*}(s-3(\ell_{2}-1)),$
since $f^{*}(3)\geq 3d-3=3(d-1)$. ∎
We abbreviate
$f^{*}_{S}\coloneqq 3(d-1)(\ell_{2}-1)+f^{*}\left(s-3(\ell_{2}-1)\right).$ (4)
The next claim concerns $G_{p}[B]$.
Claim 3.8.
Fix an obstruction. Let $B_{1},\ldots,B_{\ell_{3}}$ be the components of $G_{p}[B]$. Then,
$\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}.$
Proof.
By definition, $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq\sum_{i=1}^{\ell_{3}}f^{*}(|B_{i}|)$. Further, we have that $\sum_{i=1}^{\ell_{3}}|B_{i}|=b$, and that for each $i\in[\ell_{3}]$, $|B_{i}|\geq n/d^{C^{3}/p}$. Thus, by Remark 3.6 and by standard convexity arguments, we have that
$\sum_{i=1}^{\ell_{3}}f^{*}(|B_{i}|)$ $\geq(\ell_{3}-1)f^{*}\left(\frac{n}{d^{C^{3}/p}}\right)+\min\left\{f^{*}\left(\frac{n}{2}\right),f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}$
$\geq(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\cdot\frac{e}{C}\cdot\ln(d^{C^{3}/p})+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}$
$\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\},$
as required. ∎
Here we also abbreviate
$f^{*}_{B}\coloneqq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}+\min\left\{\frac{n}{C},f^{*}\left(b-(\ell_{3}-1)\frac{n}{d^{C^{3}/p}}\right)\right\}.$ (5)
The rest of the subsection is structured as follows. First we show that for $u$ of order $n/poly(d)$ the number of components in $B$ is at most one. Together with the trivial bound $\binom{n}{u}$ for the number of obstructions of size $u$, this suffices to show that whp there are no minimal non-trivial obstructions with $u\leq d$ (see Lemma 3.10). Then we turn to obstructions of more substantial size, particularly $d\leq u\leq\frac{n}{d^{C^{5}/p}}$; to handle such obstructions, we first establish efficient enumeration of minimal obstructions of such sizes (Lemma 3.13 and 3.14), and then, utilising these enumerations, show that whp there are no minimal obstructions with such values of $u$ in Lemma 3.15. Then, we turn to obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$. Here, the focus will turn to the set $B$. We will distinguish the cases $b\leq n/2$ and $b>n/2$. If $e(B,B^{C})$ is large, the calculation is straightforward (see Lemma 3.16 and Lemma 3.19). However, if the edge boundary of $B$ is small, the fingerprint argument from Lemma 3.5 comes into play, and allows us to resolve these cases in Lemma 3.17 and Lemma 3.20.
No minimal obstructions with $u\leq\frac{n}{d^{C^{5}/p}}$ and $\ell_{3}\geq 2$.
We begin by showing that for $u\leq\frac{n}{d^{C^{5}/p}}$ whp there are no minimal obstructions such that there are at least two large components in $B$.
@@
Whp, there are no minimal obstructions with $\ell_{3}\geq 2$ and $u\leq\frac{n}{d^{C^{5}/p}}$.
Proof.
Fix $u$, and fix a minimal obstruction $U$ with $|U|=u$. Denote by $B_{1},\ldots,B_{\ell_{3}}$ the components of $G_{p}[B]$. By Claim 3.8, we have that $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq f^{*}_{B}$. Since $G$ is $d$-regular, $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}\setminus U)\geq f^{*}_{B}-du$. Note that we may have counted some of the edges twice in the above, and therefore $\left|\bigcup_{i=1}^{\ell_{3}}E(B_{i},B_{i}^{C}\setminus U)\right|\geq(f^{*}_{B}-du)/2$. Note that none of these edges are present in $G_{p}$. There are at most $\binom{n}{u}$ ways to choose $U$, which uniquely defines the obstruction, and thus by the union bound the probability that for a fixed $u$ there is such an obstruction is at most
$\binom{n}{u}(1-p)^{(f^{*}_{B}-du)/2}$ $\leq n^{u}(1-p)^{n\ln d/(3d^{C^{3}/p})}$
$\leq n^{u}\left(\frac{\omega}{n}\right)^{(n\ln d)/(3d^{C^{3}/p+1})}$
$\leq\exp\left\{\ln n\left(u-\frac{n}{4d^{C^{3}/p+1}}\right)\right\}=o(1/n).$
where in the first inequality we used (5) together with our assumption that $\ell_{3}\geq 2$ and $u\leq\frac{n}{d^{C^{5}/p}}$, in the second inequality we used that $(1-p)^{d}=\omega/n$, the penultimate inequality follows since $\omega$ tends arbitrarily slow to infinity, and the last equality follows since we assume $u\leq\frac{n}{d^{C^{5}/p}}$. Union bound over the less than $n$ choices of $u$ completes the proof. ∎
Note that in the above, our choice for the upper-bound on the size of a component in $S$ (and thus a lower-bound on the size of a component in $B$) came into play.
No minimal non-trivial obstructions with $u\leq d$.
We now turn our attention to ‘small’ obstructions.
Lemma 3.10.
Whp, there are no minimal non-trivial obstructions with $u\leq d$.
Proof.
Fix $u$, and fix a minimal non-trivial obstruction. Note that by Lemma 3.9, we may assume that $\ell_{3}\leq 1$. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
Note that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b).$
Noting that for each $3\leq k\leq\frac{n}{d^{{C^{3}}/p}}$, we have that $f^{*}(k)>k$, we have that
$\hat{e}\geq\ell_{1}d+w(d-1)+s+f^{*}(b).$
Observe that if either $\ell_{1}+w\geq n/10$ or $s\geq n/10$, we would have that the probability of such an obstruction is at most
$\sum_{u=1}^{d}\binom{n}{u}(1-p)^{(\hat{e}-du)/2}\leq dn^{d}\cdot\left(\frac{\omega}{n}\right)^{n/30}=o(1).$
Thus, we may assume that $\ell_{1}+w\leq n/10$ and $s\leq n/10$ (and in particular, $b\neq 0$). In that case, $f^{*}(b)=f^{*}(s+\ell_{1}+w+u)$ is increasing with $s$, and thus our lower bound estimate for $\hat{e}$ is increasing with $s$. We consider two cases separately. First, suppose $s\neq 0$, and thus our estimate is minimised when $s=3$ (and $\ell_{2}=1$). Then $\ell_{1}\geq u-1$ and since our lower bound estimate is also increasing in $\ell_{1}$, our estimate is minimised when $\ell_{1}=u-1$. We then have that
$f^{*}(s+\ell_{1}+w+u)=f^{*}(2u+2)\geq(2u+2)(d-(C-1)\log_{C}(2u+2))\geq(2u+2)(d-2C\ln d).$
Then, the probability of such an obstruction is at most
$\sum_{u=1}^{d}n^{u}(1-p)^{\left(d(u-1)+(2u+2)(d-2C\ln d)-du\right)/2}\leq\sum_{u=1}^{d}n^{u}\left(\frac{\omega}{n}\right)^{u+1/3}\leq d\cdot n^{-1/4}=o(1),$
where we used that $\omega$ tends to infinity arbitrarily slowly.
Now, we turn to the case where $s=0$. In this case, since we consider non-trivial obstructions, we must have that $\ell_{1}\geq u+1$. Then, $\hat{e}\geq(u+1)d+f^{*}(2u+1)\geq(u+1)d+(2u+1)(d-2C\ln d)$. Then, the probability of such an obstruction is at most
$\sum_{u=1}^{d}n^{u}\left(1-p\right)^{((u+1)d+(2u+1)(d-2C\ln d)-du)/2}\leq\sum_{u=1}^{d}n^{u}\left(1-p\right)^{du+d-C\ln d}\leq d\cdot n^{-1/2}=o(1),$
completing the proof. ∎
Note, that in the last part of the proof it was crucial to exclude trivial obstructions.
No minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$.
We now turn our attention to obstructions of more substantial size.
The following fairly simple observation will be of use throughout the proof.
Claim 3.11.
Suppose $d\leq u\leq\frac{n}{d^{C^{5}/p}}$. Then, whp, there are no obstructions with either $\ell_{1}+w\geq n/10$ or $s\geq n/10$.
Proof.
Similar to the calculation in Lemma 3.10 note that if either $\ell_{1}+w\geq n/10$ or $s\geq n/10$, we would have that the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{n}{u}(1-p)^{(\hat{e}-du)/2}\leq\frac{n}{d^{C^{5}/p}}n^{n/d^{C^{5}/p}}\cdot\left(\frac{\omega}{n}\right)^{n/(30d)}=o(1).$
It will also be of use to rule out the existence of such minimal obstructions with $\ell\geq 1.1u$.
Lemma 3.12.
Whp, there are no minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$ and $\ell\geq 1.1u$.
Proof.
Fix $u$, and fix a minimal obstruction. Note that by Lemma 3.9, we may assume that $\ell_{3}\leq 1$. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
As before, we note that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b),$
and since for each $3\leq k\leq\frac{n}{d^{{C^{3}}/p}}$, we have that $f^{*}(k)>k$, we obtain
$\hat{e}\geq d\ell_{1}+w(d-1)+s+f^{*}(s+\ell_{1}+w+u).$
By Claim 3.11, we may assume that $\ell_{1}+w\leq n/10$ and $s\leq n/10$ (and in particular, $b\neq 0)$. In that case, $f^{*}(b)=f^{*}(s+\ell_{1}+w+u)$ is increasing with $s,\ell_{1}$ and $w$, and thus our lower bound estimate for $\hat{e}$ is increasing with $s,\ell_{1}$ and $w$. Thus, our estimate is minimised when $s=0$ (and $\ell_{2}=0$), $w=0$ and $\ell_{1}=1.1u$. We then have that
$f^{*}(s+\ell_{1}+w+u)=f^{*}(2.1u)\geq 2.1u(d-C\log_{C}(2.1u))$
Then, the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}n^{u}(1-p)^{\left(1.1du+2.1u(d-2C\ln u)-du\right)/2}\leq\sum_{u=d}^{n/d^{C^{5}/p}}n^{u}\left(\frac{\omega}{n}\right)^{1.05u}=o(1),$
where we used that $\omega$ tends to infinity arbitrarily slowly and $u\geq d$. ∎
We can now count the number of (relevant) obstructions much more efficiently.
Lemma 3.13.
Suppose $u\geq d$, $\ell\leq 1.1u$ and $\ell_{3}=1$. Then, there are at most
$\binom{n}{\frac{5}{11}u}(ed)^{\frac{6}{11}u+\ell_{1}+s}$
choices for a minimal obstruction.
Proof.
By Lemma 3.2 each vertex $v\in U$ sends edges to at least three components in $G_{p}[V_{1}\cup S\cup B]$, and by our assumptions that $\ell_{3}=1$, it neighbours at least two components in $G_{p}[V_{1}\cup S]$.
Consider now the following bipartite auxiliary graph $F$ with $V(F)=(A_{1},A_{2})$. Set $A_{1}=U$, and let $A_{2}$ be the set of components in $G_{p}[V_{1}\cup S]$. We draw an edge between two vertices $v\in A_{1}$ and $C\in A_{2}$ in this auxiliary graph $F$ if $v$ sends an edge to the component $C$ in $G_{p}$. Note that, by our assumptions, there are at least $2u$ edges in $F$. On the other hand, $|A_{2}|=\ell-1\geq u$ and thus the average degree in $F$ is at least $\frac{2u}{\ell-1}\geq\frac{20}{11}$, using that $d\leq\ell\leq 1.1u$. In a component of size at most three the average degree is at most $\frac{4}{3}$, thus at most $\frac{3}{11}$ of the vertices in $A_{1}$ lie in a component of size at most three (otherwise, the average degree would be less than $\frac{4}{3}\cdot\frac{3}{11}+2\cdot\frac{8}{11}=\frac{20}{11}$).
Going back to the original graph $G$, this remains true, i.e., at least $\frac{8}{11}u$ vertices of $U$ lie in components of size at least four in $G$. Thus, the number of choices for such an obstruction is at most
$\binom{n}{\frac{5}{11}u}(ed)^{\frac{6}{11}u+\ell_{1}+s}.$
In fact, we will use the above counting in order to provide a typical better estimate on the number of minimal obstructions. Here, we will first assume that $s=0$; we will later show that indeed the probability of an obstruction is maximized when $s=0$ (similarly to Lemma 3.12).
Lemma 3.14.
@@
Suppose $d\leq u\leq n/d^{C^{5}/p}$, $\ell\leq 1.1u$, $s=0$ and $\ell_{3}=1$. Then, whp, for every minimal obstruction, there is a $M\subseteq V$ satisfying the following,
@@
    $G[M]$ contains at most $u/(2^{d/(600C)})$ components.
@@
By Lemma 3.13, there are at most $\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}+s}$ choices for the obstruction.
In $G$ each vertex in $V_{1}$ sends $d$ edges to other components. Since $U$ is an obstruction in $G_{p}$ all the remaining edges need to lie in $E(U,V_{1})$. Suppose first that $e(U,V_{1})\leq\frac{du}{20}$. Then, we have that
$\hat{e}-e(U,V_{1})\geq d\ell_{1}+f^{*}(b)-du/20\geq d\ell_{1}-du/20.$
Thus, the probability of such an obstruction is at most
$\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}}(1-p)^{(d\ell_{1}-du/20)/2}\leq\left(\frac{11en}{5u}\right)^{5u/11}(ed)^{6u/11+\ell_{1}}\left(\frac{\omega}{n}\right)^{\ell_{1}/2-u/40}.$
Noting that the above is decreasing as $\ell_{1}$ is increasing, and that under our assumptions $\ell_{1}\geq u$, we obtain that the probability of such an obstruction is at most
$\left(\frac{11en}{5u}\right)^{5u/11}(ed)^{17u/11}\left(\frac{\omega}{n}\right)^{19u/40}\leq n^{-u/50},$
where we used that $\omega$ is a function tending arbitrarily slowly to infinity and that $d=\Theta(\ln n)$. Summing over all possible values of $u$, we obtain that whp there is no such obstruction. In what follows, we may thus assume that $e(U,V_{1})>du/20$.
Let $M=U\cup V_{1}$. We have that $U\subseteq M$ and $|M|=u+\ell_{1}\leq u+1.1u\leq 3u$, by our assumption $\ell_{1}<\ell\leq 1.1u$. 1.3 implies that for any set $M\subseteq V(G)$
@@
If $G[M]$ had at least $u/2^{\frac{d}{600C}}$ components, then by standard convexity arguments, we would have that
$e_{G}(M)$ $\leq|M|(C-1)\log_{2}\left(\frac{|M|}{u/2^{\frac{d}{600C}}}\right)$
$\leq 3u(C-1)\log_{2}\left(3\cdot 2^{\frac{d}{600C}}\right)<\frac{du}{20},$
@@
With these lemmas at hand, we are ready to show that whp there are no minimal obstructions for this range of $u$.
@@
Whp, there are no minimal obstructions with $d\leq u\leq\frac{n}{d^{C^{5}/p}}$.
@@
Assume that there is an obstruction with $\sqrt{d}\leq u\leq\frac{n}{d^{C^{5}/p}}$. By Lemma 3.12 and Lemma 3.9, we may assume that $\ell\leq 1.1u$ and $\ell_{3}\leq 1$.
Fix $u$, and fix a such a minimal non-trivial obstruction. Let $W_{1},\ldots,W_{w/2}$ be the components in $G_{p}[W]$. Let $S_{1},\ldots,S_{\ell_{2}}$ be the components in $G_{p}[S]$. Denote by $v_{1},\ldots,v_{\ell_{1}}$ the vertices of $V_{1}$.
We claim that the function
$\binom{n}{5u/11}(ed)^{\frac{6}{11}u+\ell_{1}+s}(1-p)^{(\hat{e}-du)/2}$ (6)
is decreasing as either $w$ or $s$ are increasing, and since (6) is an upper bound on the probability of an obstruction we may assume that $w=0$ and $s=\ell_{2}=0$ (and union bound over the less than $n^{2}$ choices of $s$ and $w$).
Indeed, recall that
$\hat{e}\geq d\ell_{1}+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+f^{*}(b),$
where $\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})\geq w(d-1)$, $\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})\geq f^{*}_{S}$ and $f^{*}(b)\geq f^{*}(\ell_{1}+w+s+u)$. By 3.11 we may assume that $s<n/10$ and $w<n/10$. Then $\ell_{1}+w+s+u\leq n/4$ (recalling that $\ell_{1}\leq\ell\leq 1.1u$) and $f^{*}(b)=f^{*}(\ell_{1}+w+s+u)$ is increasing with $w$ and $s$.
Thus, in order to show that (6) is decreasing as $s$ is increasing, it suffices to observe that $(ed)^{s}(1-p)^{f^{*}_{S}}$ is decreasing when $s$ is increasing (recalling that $f^{*}_{S}\coloneqq 3(d-1)(\ell_{2}-1)+f^{*}(s-3(\ell_{2}-1))$ from (4)). Indeed, increasing the value of $s$ will not force us to increase the value of $\ell_{1}$ (in fact, it may do the opposite), and as we noted $f^{*}(\ell_{1}+w+s+u)$ is increasing with $s$.
Similarly, $(ed)^{w/2}(1-p)^{w(d-1)}$ is decreasing when $w$ is increasing. Thus, since we observed that $f^{*}(\ell_{1}+w+s+u)$ is increasing with $w$, (6) is decreasing as $w$ is increasing.
@@
Assuming $w=s=0$, by Lemma 3.14, we may assume that there exists $M$ such that $|M|\leq 3u$, $U\subseteq M$ and there are at most $u/2^{\frac{d}{(600C)}}\eqqcolon r$ components in $G[M]$. We have at most $n^{r}$ ways to choose the sizes of these components. Thus, by Lemma 1.4, there are at most $\binom{3u}{u}n^{r}(ed)^{3u}$ ways to choose $U$. Thus, the probability of having such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{3u}{u}n^{r}(ed)^{3u}(1-p)^{(\hat{e}-du)/2},$
where $\hat{e}\geq d\ell_{1}+f^{*}(b)=d\ell_{1}+f^{*}(2u)\geq du+\frac{2eu}{C}\cdot\ln(n/2u)$. We thus have that
$\sum_{u=d}^{n/d^{C^{5}/p}}\binom{3u}{u}n^{r}(ed)^{3u}(1-p)^{(\hat{e}-du)/2}$
$\leq$ $\sum_{u=d}^{n/d^{C^{5}/p}}(3e)^{u}n^{u/2^{\frac{d}{(600C)}}}(ed)^{3u}\exp\left\{-p\cdot\frac{eu}{C}\cdot\ln(n/2u)\right\}$
$\leq$ $\sum_{u=d}^{n/d^{C^{5}/p}}\exp\left\{u\left(\ln(3e)+\frac{\ln n}{2^{d/(600C)}}+3\ln(ed)-\frac{ep}{C}\ln(n/2u)\right)\right\}.$
@@
Recalling that $\ln n=\Theta(d)$ and using that $u\leq n/d^{C^{5/p}}$ the probability of such an obstruction is at most
$\sum_{u=d}^{n/d^{C^{5}/p}}\exp\left\{-uC^{3}\ln d\right\}=o(1/n^{2}),$